OpenStudy (anonymous):

an inventor bought 600 shares of stock, some at $6.25 per share and some at 1.00 per share. If the total cost was $1387.50, how many shares of each stock did the investor buy? (round to two decimal places if necessary)

OpenStudy (whpalmer4):

call the shares bought at $6.25 \(x\) and the shares bought at $1 \(y\) \[x+y = 600\]\[6.25x+1y=1387.50\]Solve the system of equations to find \(x\) and \(y\). Do you know how?

OpenStudy (anonymous):

I am not sure because I dont know where to start. Could you walk me through it please?

OpenStudy (***[isuru]***):

do u mean u need help to solve the system of equation ?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Pick one of the variables in the first equation and solve for it. You'll get either \[x = 600-y\]or\[y=600-x\]Now replace the variable on the left side with the expression on the right side wherever you see that variable in the second equation. That gives you an equation in terms of one variable, which you should be able to solve without assistance. Find the value of that variable, then plug it into the first equation to find the value of the other variable.

OpenStudy (whpalmer4):

That's called the substitution method. Another method is called elimination, but it's time for me to call it a night, maybe Isuru will explain it to you.

OpenStudy (anonymous):

okay then thank you

OpenStudy (whpalmer4):

were you able to solve the problem?

OpenStudy (anonymous):

no I wasnt. Im confused of where to start

OpenStudy (whpalmer4):

Pick a letter, \(x\) or \(y\)

OpenStudy (whpalmer4):

Well?

OpenStudy (anonymous):

I picked y and I got the answer x=162 and y=-102?

OpenStudy (whpalmer4):

\[x+y = 600\]Solve for \(y\):\[x-x+y=600-x\]\[y=600-x\]Did you get that?

OpenStudy (anonymous):

yeah I did.

OpenStudy (whpalmer4):

Okay. What did you get when you substituted that into the other equation?

OpenStudy (anonymous):

I got 3750-6.25x=1387.50. I got -378?

OpenStudy (whpalmer4):

How the heck did you get that? \[6.25x + 1y = 1387.50\]Right?

OpenStudy (whpalmer4):

Now replace \(y\) with \((600-x)\), what do you get?

OpenStudy (whpalmer4):

\[6.25x + 1(600-x) = 1387.50\]That's the result of the substitution of \(y = 600-x\) Solve that for \(x\), please.

OpenStudy (anonymous):

I got 150?

OpenStudy (anonymous):

6.25 per share =x=150

OpenStudy (whpalmer4):

How many shares at the $1 price?

OpenStudy (anonymous):

I got y=450?

OpenStudy (whpalmer4):

Good. Let's check the result: 150 shares @ 6.25 + 450 @ 1 150*6.25 + 450*1 = 937.5 + 450 = 1387.50 that part checks 150 + 450 = 600 that part also checks It's important to test the solution in all of the original equations, because it is possible to construct erroneous solutions which will satisfy some of the original equations, but not all. For example, 500 of the cheap stock and 100 of the expensive gives you 600 shares, but won't satisfy the investment pricing equation. We could also make up a set of numbers that satisfies the pricing equation, but doesn't satisfy the share count.

OpenStudy (whpalmer4):

I suppose I could show you the elimination technique if you are interested.

OpenStudy (anonymous):

Okay then. The more I learn, the better

OpenStudy (whpalmer4):

Good answer! :-) So with elimination, we want to find a set of operations we can do to combine the equations in such a fashion that we eliminate variables until we are left with an equation in only one variable to solve. Then we use that variable to find the values of the other variable(s), much like we did in substitution. \[x+y=600\]\[6.25x+1y=1387.50\] We would like to multiply one or both of the equations by numbers chosen so that we end up with one of the variables having coefficients which are equal in magnitude but opposite in sign. For example, in this problem, we've got \(y\) in the first equation and \(1y\) in the second. If we multiply one of the equations by -1, we'll have a \(1y\) and a \(-1y\), and if we add the equations together, that will give us \(0y\), leaving us an equation just in terms of \(x\). Does that make sense so far?

OpenStudy (anonymous):

yes it does.

OpenStudy (whpalmer4):

Multiplying the whole equation by the same number doesn't change the solution at all because we are doing the same thing to both sides of the equation. 2x=4 has the same solution as 2*2x = 4*2, right?

OpenStudy (anonymous):

okay

OpenStudy (whpalmer4):

Okay, so let's do just that: \[-1*x -1*y = -1*600\]\[6.25x+1y=1387.50\] ------------------------------- \[6.25x-1x + 1y-1y = 1387.50-600\]\[5.25x=787.50\]Solve that for \(x\), then plug the value into either of the original equations to find \(y\).

OpenStudy (whpalmer4):

Here it was easy because we already had a variable where the coefficients were equal. We could have just subtracted one equation from the other instead of multiplying by -1 and adding. Sometimes you'll have a more complicated setup, like this: \[2x+3y=6\]\[3x+4y=0\] Here I would multiply the first equation by -3, and the second by 2: \[-3*2x-3*3y=-3*6\]\[2*3x+2*4y=2*0\] \[-6x-9y=-18\]\[6x+8y=0\]--------------\[0x-y=-18\]\[y=18\]\[2x+3(18)=6\]\[2x+54=6\]\[x=-24\]

OpenStudy (whpalmer4):

It's a lot like making a common denominator when adding fractions: \[\frac{1}{2} +\frac{1}{3} = \frac{1}{2}*\frac{3}{3} + \frac{1}{3}*\frac{2}{2}\]

OpenStudy (anonymous):

okay that makes sense.

OpenStudy (whpalmer4):

just multiply by the coefficient from the other equation, changing the sign if necessary. sometimes you get lucky and the coefficient in one equation is a multiple of that in the other equation, in which case you only have to multiply one equation by something.

OpenStudy (whpalmer4):

so that's elimination. you can do it on larger problems, like 3 equations in 3 unknowns. you just have to repeat the process to get rid of each additional variable. if you have 3 eqs in 3 unks, you would first eliminate a variable from the first pair of equations (1st and 2nd), and then the same variable from the second pair of equations (2nd and 3rd). Now you have two new equations in only two unknowns, and you solve them like we did here. With those solved, you go back and find the third variable's value.

OpenStudy (anonymous):

Thank You so much for helping me understand all of this. you are awesome!!

OpenStudy (whpalmer4):

You can do larger systems with substitution as well, but the algebra gets messier.

OpenStudy (whpalmer4):

Yet another way to solve a system of equations like this is by graphing. If the two lines intersect, the coordinates of the point at which they intersect is the solution. Here's a graph showing that for this problem.

OpenStudy (whpalmer4):

If they don't intersect (because they are parallel), there is no solution. If the lines are identical, there are an infinite number of solutions. When you solve a system where there is no solution, you'll end up with something crazy like \[4=0\]A false result like that means no solution, because the lines are parallel. If you instead get something like \[0=0\]that means the lines are coincident (the same) and there are infinitely many solutions. Any point on the line represents a solution in this case.